Calculte the no of total neutron of 25 gm ammoniN=14 h=1

Grade:11

1 Answers

Aditya yadav
6 years ago
molecular mass of ammonium=14*(1)+4*(1)=18au
1 mole of NH4 = 18gm of NH4
therefore 1/18 mole = 1 gram ammonium
now, for 25 grams = (1/18)*25
=1.38 moles
=(1.38)*6.022*1023
=8.31 * 1023 molecules of ammonium
now in molecule there are 7 neatrons that too of only nitrogen as hydrogen has no neutrons.
so, in 8.31 * 1023 molecules of ammonium there will be 7*8.31 * 1023 neutrons
which is equal to 5.81*1023 neutrons (answer)